A concrete mixture needs 660 kg dry sand and 1240 kg dry stone. It is known that the water


content of the existing sand is 4% and that of the stone is 1%. Calculate the amount of wet


sand and wet stone respectively.

1
Expert's answer
2022-02-03T20:27:02-0500

% moisture content= ((Weight of wet sand (M)- weight of dry sand (D))/ Weight of dry sand (D)

4%= (M-660kg)/660kg

0.04*660kg= M-660kg

26.4kg=M-660kg

M=660kg+26.4kg

M= 686.4kg

Weight of wet sand= 686.4kg



% moisture content= ((Weight of wet stone (S)- weight of dry stone (D))/ Weight of dry stone (D)

1%=(S-1240kg)/1240kg

0.01*1240kg= S-1240kg

12.4kg=S-1240

S=1252.4kg

weight of wet stone= 1252.4kg